At 2:15 o'clock the hour and the minute hands of a clock form an angle of ...degrees
Answer by MissionPossible(18)

The Circle has 360 degree
the clock has 12 point of 12 hour.
Then each point would has 360/12 = 30 degree.
At 2 o'clock it would be at 30x2 = 60 degree
At 2.15 the minute hand would be at 90 degree.
and the hour hand would move from 60 degree to 60 + (15/60 * 30) degree
(15/60 * 30) is explained by
for each hour the hour hand would move 30 degree
for 15 minute it would move 15/60 * 30 = 7.5
Then at 2.15 o'clock the hour hand would be at 67.5 degree and the minute hand would be at 90 degree.
Then they would form an angle of 90 - 67.5 = 22.5 degree